1.A company is planning the launch of a new product. It must choose one product from three available projects and has estimated the following data.
Product A Product B Product C
Annual Sales 680 900 1200
Unit Cost 680 900 1200
Fixed Cost 200000 350000 500000
Product life 3 years 5years 8years
Selling price 760 1000 1290
a.Appraise and recommend a product.
b.Write the total project cost of the company if all three products are launched?
c.Point out the environmental factors you would recommend the company in making decisions regarding the new product?
1
Expert's answer
2019-05-08T10:10:54-0400

profit from the sale of 1 piece of product : A= 760-680 = 80, B= 1000-900 = 100 , C= 1290-1200 = 90

annual sales profit A= 80*680 = 54400 , B=100*900 = 90000 , C= 90*1200 = 108000

profit from sales for the entire period A= 54400*3 =163200 , B= 90000*5 = 450000 , C= 108000*8 = 864000

total project cost = 680*680*3 + 900*900*5 + 1200*1200*8 = 1387200 + 4050000 + 11520000 = 16957200

I recommend running project C.

Of the environmental factors - I advise you to analyze competitors, to conduct a marketing research - a survey